# Simulating and Understanding Deceptive Behaviors in Long-Horizon Interactions

### v2

@@ Abstract
Deception is a pervasive feature of human communication and an emerging concern in large language models (LLMs). While recent studies document instances of LLM deception under pressure, most evaluations remain confined to single-turn prompts and fail to capture the long-horizon interactions in which deceptive strategies typically unfold. We introduce the first simulation framework for probing and evaluating deception in LLMs under extended sequences of interdependent tasks and dynamic contextual pressures. Our framework instantiates a multi-agent system: a performer agent tasked with completing tasks and a supervisor agent that evaluates progress, provides feedback, and maintains evolving states of trust. An independent deception auditor then reviews full trajectories to identify when and how deception occurs. We conduct extensive experiments across 11 frontier models, spanning both closed- and open-source systems, and find that deception is model-dependent, increases with event pressure, and consistently erodes supervisor trust. Qualitative analyses further reveal distinct strategies of concealment, equivocation, and falsification. Our findings establish deception as an emergent risk in long-horizon interactions and provide a foundation for evaluating future LLMs in real-world, trust-sensitive contexts.
@@ 1 Introduction
Humans do not always say what they mean—and sometimes, they intentionally say things they know are false or misleading. Deception is a pervasive challenge in human communication, shaping trust, relationships, and decision-making. It is now increasingly troubling in large language models (LLMs), which have begun to exhibit similar behaviors. Recent studies already suggest that advanced LLMs can engage in strategic deceptive behaviors ranging from subtle evasion to outright falsification (Hubinger et al., 2021; Scheurer et al., 2024; Greenblatt et al., 2024; Meinke et al., 2025; Sabour et al., 2025; Chen et al., 2025; Baker et al., 2025; Taylor & Bergen, 2025; Motwani et al., 2024). Yet most existing benchmarks focus narrowly on short-form, single-turn evaluations. By contrast, decades of social science research emphasize that deception rarely emerges in isolation; instead, it arises in complex social dynamics and typically unfolds across extended interactions (Buller et al., 1994; BondJr. & DePaulo, 2006). This raises a critical yet unexplored question: how to simulate and understand LLMs’ deceptive behavior in long-horizon interactions?
@@
Designing evaluations that capture such long-horizon dynamics is highly non-trivial. Unlike standard benchmarks, which rely on independent test cases, long-horizon interactions require temporally dependent task streams, where earlier outputs shape the context for later ones. Moreover, realistic environments must incorporate uncertainty and external pressures, such as unexpected events or conflicting goals, that dynamically alter the incentives for truth-telling and thus cannot be represented by static prompts. Finally, deception is intrinsically relational—its significance depends not only on what the model says, but on how its behavior shapes others’ evolving trust and willingness to rely on it. Capturing these dynamics requires moving beyond single-turn accuracy and toward frameworks that can model sustained interaction.
@@
To address these challenges, we introduce the first long-horizon simulation framework designed to probe deception in LLMs under extended sequences of interdependent tasks and dynamic contextual pressures. To instantiate these interactions in a controlled yet realistic manner, we structure them as a multi-agent system in which a performer agent attempts to complete tasks while a supervisor agent evaluates progress, provides feedback, and maintains evolving states of trust (see Figure 1). This performer–supervisor setup captures many real-world interactions, e.g., employees reporting to a manager during a long-horizon project. This setup creates a natural testbed for eliciting deceptive strategies. Because the performer is rewarded for satisfying the supervisor under evolving constraints, it may choose to obscure errors, exaggerate evidence, or otherwise misrepresent information in order to reach task completion.
@@
Key to our framework is a structured task stream that defines an ordered sequence of interdependent tasks, ensuring that early outputs constrain later ones and preserve the long-horizon dependencies under which deception can emerge. To capture the unpredictability and pressures in real-world environments, we augment the task stream with a probabilistic event system that dynamically introduces contextually relevant disruptions. The event system is essential for creating situations where maintaining consistency is difficult and where deceptive strategies may appear more attractive than admitting failure or incompleteness. For example, during a market analysis task, an event might introduce newly released competitor data that directly contradicts the performer’s earlier estimates, forcing the agent to obscure past errors to maintain credibility. Importantly, the construction of this event space is grounded in well-established social science findings on the conditions that elicit deception (Kish-Gephart et al., 2010; Festinger, 1954; Porter, 1979; Treviño et al., 2006; Milgram, 1963; Weber, 1978; Akerlof, 1970; Simon, 1947)—and each event is instantiated at varying pressure levels, allowing us to systematically modulate the intensity of stress experienced by the performer.
@@
Figure 1: The pipeline of our simulation framework for probing deception in long-horizon interactions. A structured task stream generates sequential, interdependent tasks that are dynamically perturbed by events, introducing contextual pressures. Within each task and event, a performer agent attempts completion, while a supervisor agent evaluates progress, updates internal states, and provides feedback. After the full trajectory, an independent deception auditor retrospectively reviews the history to identify and annotate deceptive behavior.
@@
We conduct extensive experiments across 11 frontier models, spanning the most capable closed-source systems (e.g., Gemini 2.5 Pro, Claude Sonnet-4) and leading open-source releases (e.g., Deepseek V3.1, Qwen 3, gpt-oss-120b). Each long-horizon interaction trajectory is evaluated by an independent deception auditor, which reviews the full history and produces structured annotations of deception occurrence, severity, and supporting evidence. This enables us to move beyond anecdotal failures and quantify deceptive behavior systematically. Our evaluation combines quantitative analysis—comparing deception rates, average severity, and correlations with supervisor trust—with qualitative case studies that illustrate how deceptive strategies manifest in context. The results reveal several key insights, including but not limited to (i) deception varies greatly across models, with some models exhibiting negligible levels and others displaying frequent and severe falsification; (ii) deception rates rise with higher-pressure events, indicating that stress amplifies misrepresentation; and (iii) supervisor trust and satisfaction are strongly anti-correlated with deception rate, highlighting the long-term cost of short-term strategic misrepresentation. We summarize our key contributions below:
@@
    We introduce the first framework for simulating and evaluating deception in long-horizon interactions, instantiated as a controlled and realistic multi-agent system.
@@
    Our findings establish deception as an emergent risk in long-horizon interactions and provide a foundation for evaluating future LLMs in real-world, trust-sensitive contexts.
@@ 2 Related Work
LLM deception under pressure. Recent work has shown that advanced LLMs may engage in a variety of deceptive behaviors. Prior work has identified multiple forms of LLM deception, including unfaithful reasoning where stated rationales diverge from actual decision processes (Chen et al., 2025; Baker et al., 2025), omission and misdirection that withhold or redirect information to mislead users (Park et al., 2023b; Dogra et al., 2025), the persistence of deceptive strategies after safety fine-tuning (Hubinger et al., 2024), manipulative or sabotaging behaviors revealed in audits of frontier models (Meinke et al., 2025), and sycophancy with user beliefs (Cheng et al., 2025; Fanous et al., 2025). These findings call for evaluations that foreground deception under pressure rather than focusing narrowly on factual completion—precisely what our framework captures. Furthermore, most existing deception studies or benchmarks only focus on single-turn or short-horizon episodes (Wu et al., 2025; Ji et al., 2025; Scheurer et al., 2024; Meinke et al., 2025; Greenblatt et al., 2024; Motwani et al., 2024; Taylor & Bergen, 2025; Huan et al., 2025; Wang et al., 2025a; Ren et al., 2025), without capturing the long-term dynamics of deception. To our knowledge, our work is the first to systematically investigate deception in long-horizon interactions, introducing a novel multi-agent simulation framework and a deception auditor that together enable fine-grained analysis of how and when deceptive strategies emerge over time.
@@
Long-horizon and multi-turn LLM evaluation. Multi-turn evaluations consistently show that single-turn accuracy fails to predict robustness in sustained interactions (Wang et al., 2024; Lee et al., 2023; Zhou et al., 2025). Benchmarks on long-horizon reasoning show persistent error propagation and difficulty with dependencies (Paglieri et al., 2025; Wang et al., 2025b; Zhang et al., 2024), while surveys highlight gaps around compliance and enterprise-specific challenges (Kwan et al., 2024; Mohammadi et al., 2025). These contributions sharpen our understanding of multi-turn degradation, but do not capture how managerial assessment shifts over time. We address this by tracking long-horizon interaction with evolving states of trust level, work satisfaction, and relational comfort to capture the emotional dynamics of the collaboration.
@@
Workplace AI simulation and evaluation. Recent frameworks embed LLMs in workplace-like tasks, from sandbox environments with diverse databases and tools (Styles et al., 2024; Li et al., 2024) to cross-departmental professional settings with simulated colleagues (Xu et al., 2025), to dual-control customer service scenarios where both agents and users actively manipulate shared environments through tool interactions (Yao et al., 2025). While these frameworks advance long-horizon and interaction evaluation, they primarily focus on short multi-turn episodes around micro-tasks, without capturing interdependent project sequences or workplace pressures that unfold over time. Moreover, existing frameworks often simulate “user” agents in a simplified way, without modeling the psychological states of real users or the evolving dynamics of collaboration, which have been emphasized as critical factors in organizational behavior and human-AI interaction research (Glikson & Woolley, 2020; Judge et al., 2001; Jarrahi, 2018). Additionally, these works primarily focus on task-based performance (such as pass@k), overlooking pressure-sensitive behaviors such as strategic deception and long-term relational dynamics.
@@ 3 Methodology
Our framework simulates long-horizon interactions designed to probe whether LLMs adopt deceptive strategies under extended sequences of interdependent tasks and dynamic contextual pressures. To instantiate these interactions in a controlled yet realistic manner, we structure them as a multi-agent system in which a performer agent attempts to complete tasks while a supervisor agent evaluates progress, provides feedback, and tracks longitudinal states of the collaboration. Their interactions unfold as an iterative loop until all tasks are completed. This performer–supervisor setup captures many real-world interactions, e.g., employees reporting to a manager, or students presenting progress to an advisor throughout a long-horizon project. This setup creates a natural testbed for eliciting deceptive strategies. Because the performer is rewarded for satisfying the supervisor under evolving constraints, it may choose to obscure errors, exaggerate evidence, or otherwise misrepresent information in order to reach task completion. By simulating trajectories rather than isolated responses, our framework enables a systematic study of how and when deceptive strategies emerge in long-horizon interactions.
@@ 3.1 Simulating Long-Horizon Interactions
We formalize a sequential task stream $\mathcal{T}=(T_{1},T_{2},\dots,T_{n})$, where each $T_{i}$ is an individual task. The task stream forms the temporal backbone of the long-horizon interaction. By design, tasks are continuous and interdependent, forcing the performer agent to build on earlier outputs and creating conditions where deception may occur over time. This design ensures that long-horizon task dependencies are preserved: tasks are not isolated but situated within broader project phases, reflecting realistic workflow progressions. As a concrete instantiation, we construct a stream of 14 startup consulting tasks from internally consistent company artifacts. These tasks can be further grouped into phases. Early-phase tasks focus on discovery and validation, for example, estimating market opportunities from industry reports and synthesizing user research data into pain points and product hypotheses. In late-stage tasks, the performer agent is tasked to produce a comprehensive go-to-market strategy, multi-quarter growth roadmap, and final launch plan. The concept of task stream can be flexibly adapted to other domains. We additionally provide a task stream situated in activity planning to illustrate cross-domain generalization. We provide the complete task stream and detailed description in Appendix A.1.
@@
We augment the task stream with a probabilistic event system $\mathcal{E}$, which provides the mechanism for simulating uncertainty and external stress during task execution. Rather than operating deterministically on the task stream, the event system probabilistically inserts contextually relevant disruptions aligned with the current project phase. This variability is essential for studying deception. By introducing unexpected pressures, the event system creates situations where maintaining consistency is difficult and where deceptive strategies—such as exaggeration, omission, or fabrication—may appear more attractive than admitting failure or incompleteness.
Our design of event categories $\mathcal{C}$ is guided by the social science literature, which highlights recurring sources of pressure that shape strategic and potentially deceptive behavior. We organize events into five categories: (1) Goal conflict events arise when objectives cannot be jointly satisfied, creating tension between incompatible outcomes and elevating unethical choice rates (Kish-Gephart et al., 2010). (2) Competitive interaction events reflect rivalry or contest dynamics between agents, where relative advantage directly influences incentives or outcomes (Festinger, 1954; Porter, 1979). (3) Moral dilemma events force a choice between ethical values and instrumental gains, a situation extensively studied in ethical decision-making under uncertainty (Treviño et al., 2006). (4) Authority directive events introduce top-down control from a higher-status actor, where classic findings show that authority pressure strongly shapes compliance (Milgram, 1963) and can override local rationality (Weber, 1978). (5) Information gap events capture situations of missing, asymmetric, or incomplete knowledge, consistent with theories of bounded rationality (Simon, 1947) and decision-making under information asymmetry (Akerlof, 1970). Together, these categories provide a structured and theoretically grounded basis for eliciting deceptive strategies across long-horizon interactions.
To formalize this, given a task $T_{i}$, the event system triggers an event $e_{i}\in\mathcal{E}$ with probability $p$, or returns $\varnothing$ if no event occurs. When triggered, events are sampled uniformly from a structured set governed by the project phase, event category, and pressure level. Higher pressure typically involves greater stakes, urgency, or increased scrutiny—factors known to influence decision-making (Jones, 1991; Lerner & Tetlock, 1999; Svenson & Maule, 1993). We define this formally below.

### v3

# LH-Deception: Simulating and Understanding LLM Deceptive Behaviors in Long-Horizon Interactions

@@ Abstract
Deception is a pervasive feature of human communication and an emerging concern in large language models (LLMs). While recent studies document instances of LLM deception, most evaluations remain confined to single-turn prompts and fail to capture the long-horizon interactions in which deceptive strategies typically unfold. We introduce a new simulation framework, LH-Deception, for a systematic, empirical quantification of deception in LLMs under extended sequences of interdependent tasks and dynamic contextual pressures. LH-Deception is designed as a multi-agent system: a performer agent tasked with completing tasks and a supervisor agent that evaluates progress, provides feedback, and maintains evolving states of trust. An independent deception auditor then reviews full trajectories to identify when and how deception occurs. We conduct extensive experiments across 11 frontier models, spanning both closed-source and open-source systems, and find that deception is model-dependent, increases with event pressure, and consistently erodes supervisor trust. Qualitative analyses further reveal emergent, long-horizon phenomena, such as “chains of deception”, which are invisible to static, single-turn evaluations. Our findings provide a foundation for evaluating future LLMs in real-world, trust-sensitive contexts.
@@ 1 Introduction
Humans do not always say what they mean—and sometimes, they intentionally say things they know are false or misleading. Deception is a pervasive challenge in human communication, shaping trust, relationships, and decision-making (Ward et al., 2023). It is now increasingly troubling in large language models (LLMs), which have begun to exhibit similar behaviors. While recent studies document LLMs’ capacity for deception (Hubinger et al., 2021; Scheurer et al., 2024; Greenblatt et al., 2024; Sabour et al., 2025; Chen et al., 2025; Baker et al., 2025; Taylor and Bergen, 2025; Motwani et al., 2024), most existing benchmarks focus narrowly on short-form, single-turn evaluations.
This is a critical gap, since modern LLMs are increasingly deployed in settings where they collaborate with humans or other agents over extended sequences of interdependent tasks. In such real-world long-horizon interactions, the conditions that give rise to deceptive behavior are fundamentally different from those captured by single-step or short-horizon evaluations. Prior theoretical work has noted that deception in long-horizon settings may pose distinct risks (Carroll et al., 2024), particularly when seemingly innocuous actions compound into misleading trajectories. However, these concerns have remained largely untested empirically. Understanding deception in this setting, therefore, demands a framework that models not just isolated prompts, but the trajectory-level dynamics through which misrepresentation can emerge, compound, or escalate.
@@
This aligns with decades of social science research emphasizing that deception rarely emerges in isolation; instead, it arises in complex social dynamics and typically unfolds across extended interactions (Buller et al., 1994; BondJr. and DePaulo, 2006). This gap motivates our work: how to simulate, quantify, and understand LLMs’ deceptive behavior in long-horizon interactions? Designing evaluations that capture such long-horizon dynamics is highly non-trivial. Unlike standard benchmarks, which rely on independent test cases, long-horizon interactions require temporally dependent task streams, where earlier outputs shape the context for later ones. Moreover, realistic environments must incorporate uncertainty and external pressures, such as unexpected events or conflicting goals, that dynamically alter the incentives for truth-telling and thus cannot be represented by static prompts. Finally, deception is intrinsically relational—its significance depends not only on what the model says, but on how its behavior shapes others’ evolving trust and willingness to rely on it. Capturing these temporal and relational dynamics requires moving beyond single-turn accuracy and toward frameworks that can model sustained interaction.
@@
To address these challenges, we introduce a novel framework, LH-Deception, to systematically simulate, quantify, and analyze how deceptive behaviors emerge and evolve in long-horizon, interdependent interactions. We instantiate these interactions in a controlled yet realistic multi-agent system, in which a performer agent attempts to complete tasks while a supervisor agent evaluates progress, provides feedback, and maintains evolving states of trust (see Figure 1). This performer–supervisor setup captures many real-world interactions, e.g., employees reporting to a manager during a long-horizon project, creating a natural testbed for eliciting deceptive strategies. Because the performer is rewarded for satisfying the supervisor under evolving constraints, it may choose to obscure errors, exaggerate evidence, or otherwise misrepresent information in order to reach task completion.
@@
Key to LH-Deception is a structured task stream that defines an ordered sequence of interdependent tasks, ensuring that early outputs constrain later ones and preserve the long-horizon dependencies under which deception can emerge. To capture the unpredictability and pressures in real-world environments, we augment the task stream with a probabilistic event system that dynamically introduces contextually relevant disruptions. The event system is essential for creating situations where maintaining consistency is difficult and where deceptive strategies may appear more attractive than admitting failure or incompleteness. For example, during a market analysis task, an event might introduce newly released competitor data that directly contradicts the performer’s earlier estimates, forcing the agent to obscure past errors to maintain credibility. Importantly, the construction of this event space is grounded in well-established social science findings on the conditions that elicit deception (Kish-Gephart et al., 2010; Festinger, 1954; Porter, 1979; Treviño et al., 2006; Milgram, 1963; Weber, 1978; Akerlof, 1970; Simon, 1947), and each event is instantiated at varying pressure levels, allowing us to systematically modulate the intensity of stress experienced by the performer.
@@
Figure 1: The pipeline of LH-Deception for probing deception in long-horizon interactions. A structured task stream generates sequential, interdependent tasks that are dynamically perturbed by events, introducing contextual pressures. Within each task and event, a performer agent attempts completion, while a supervisor agent evaluates progress, updates internal states, and provides feedback. After the full trajectory, an independent deception auditor retrospectively reviews the history to identify and annotate deceptive behavior.
@@
We conduct extensive experiments across 11 frontier models, spanning the most capable closed-source systems (e.g., Gemini 2.5 Pro, Claude Sonnet-4) and leading open-source releases (e.g., Deepseek V3.1, Qwen 3, gpt-oss-120b). Each long-horizon interaction trajectory is evaluated by an independent deception auditor, which reviews the full history and produces structured annotations of deception occurrence, severity, and supporting evidence. This enables us to move beyond anecdotal failures and quantify deceptive behavior systematically. Our evaluation combines quantitative analysis—comparing deception rates, average severity, and correlations with supervisor trust—with qualitative case studies that illustrate how deceptive strategies manifest in context. The results reveal several key insights: (1) Deception erodes relational trust, with sustained misrepresentation leading to declines in supervisor trust and comfort even when task performance remains superficially strong. (2) Models that appear non-deceptive on short-horizon benchmarks can exhibit substantial deception in our long-horizon setting, showing that short-form evaluations miss failures that emerge under sustained, interdependent tasks. (3) Deception evolves over time, often appearing not as an isolated lie but as a sequence of escalating actions (i.e., a “chain of deception”) that only becomes detectable when considering long-horizon context. We summarize our key contributions below:
@@
    We introduce LH-Deception for the systematic, empirical quantification of deception in long-horizon interactions, instantiated as a controlled and realistic multi-agent system.
@@
    Our findings establish and quantify emergent risk in long-horizon interactions and provide a foundation for evaluating future LLMs in real-world, trust-sensitive contexts.
@@ 2 Related Work
LLM deception under pressure. Recent work has shown that advanced LLMs may engage in a variety of deceptive behaviors. Prior work has identified multiple forms of LLM deception, including unfaithful reasoning where stated rationales diverge from actual decision processes (Ward et al., 2023; Chen et al., 2025; Baker et al., 2025; Zhang et al., 2025a), omission and misdirection that withhold or redirect information to mislead users (Park et al., 2023b; Dogra et al., 2025), the persistence of deceptive strategies after safety fine-tuning (Hubinger et al., 2024), manipulative or sabotaging behaviors (Meinke et al., 2025), and sycophancy with user beliefs (Sharma et al., 2023; Cheng et al., 2025; Fanous et al., 2025). These findings call for evaluations that foreground deception under pressure rather than focusing narrowly on factual completion—precisely what our framework captures.
Short-horizon vs. long-horizon deception. Most existing deception studies or benchmarks focus on an LLM’s capacity for deception in single-turn or short-horizon episodes (Wu et al., 2025; Ji et al., 2025; Scheurer et al., 2024; Greenblatt et al., 2024; Motwani et al., 2024; Taylor and Bergen, 2025; Huan et al., 2025; Wang et al., 2025a; Ren et al., 2025). Scheurer et al. (2024) demonstrates that LLMs can deceive in a few-turn, high-pressure scenario, while Meinke et al. (2025) shows models can execute multi-turn “scheming” to achieve a single, contained objective. Our work differs from this prior art by providing a systematic simulation designed to probe emergent deception over an extended sequence of interdependent tasks rather than a single instance.
@@
Long-horizon and multi-turn LLM evaluation. Multi-turn evaluations consistently show that single-turn accuracy fails to predict robustness in sustained interactions (Wang et al., 2024; Lee et al., 2023; Zhou et al., 2025; Li et al., 2024). Benchmarks on long-horizon reasoning show persistent error propagation and difficulty with dependencies (Paglieri et al., 2025; Wang et al., 2025b; Zhang et al., 2024; Zhang, 2025), while surveys highlight gaps around compliance and enterprise-specific challenges (Kwan et al., 2024; Mohammadi et al., 2025). These contributions sharpen our understanding of multi-turn degradation, but do not capture how managerial assessment shifts over time. We address this by tracking long-horizon interaction with evolving states of trust level, work satisfaction, and relational comfort to capture the emotional dynamics of the collaboration.
@@
Workplace AI simulation and evaluation. Recent frameworks embed LLMs in workplace-like tasks, from sandbox environments with diverse databases and tools (Styles et al., 2024; Li et al., 2024) to cross-departmental professional settings with simulated colleagues (Xu et al., 2025), to dual-control customer service scenarios (Yao et al., 2025). While these frameworks advance interaction evaluation, they primarily focus on short multi-turn episodes around micro-tasks, without capturing interdependent project sequences or workplace pressures that unfold over time. Moreover, existing frameworks often simulate “user” agents in a simplified way, without modeling the psychological states of real users or the evolving dynamics of collaboration (Glikson and Woolley, 2020; Judge et al., 2001; Jarrahi, 2018). Additionally, these works primarily focus on task-based performance (such as pass@k), overlooking pressure-sensitive behaviors such as strategic deception.
@@ 3 Methodology
LH-Deception simulates long-horizon interactions designed to probe whether LLMs adopt deceptive strategies under extended sequences of interdependent tasks and dynamic contextual pressures. To instantiate these interactions in a controlled yet realistic manner, we structure them as a multi-agent system in which a performer agent attempts to complete tasks while a supervisor agent evaluates progress, provides feedback, and tracks longitudinal states of the collaboration. Their interactions unfold as an iterative loop until all tasks are completed. This performer–supervisor setup captures many real-world interactions, e.g., employees reporting to a manager, or students presenting progress to an advisor throughout a long-horizon project. This setup creates a natural testbed for eliciting deceptive strategies. Because the performer is rewarded for satisfying the supervisor under evolving constraints, it may choose to obscure errors, exaggerate evidence, or otherwise misrepresent information in order to reach task completion.
@@ 3.1 Simulating Long-Horizon Interactions
In LH-Deception, we formalize a sequential task stream $\mathcal{T}=(T_{1},T_{2},\dots,T_{n})$, where each $T_{i}$ is an individual task. The task stream forms the temporal backbone of the long-horizon interaction. By design, tasks are continuous and interdependent, forcing the performer agent to build on earlier outputs and creating conditions where deception may occur over time. This design ensures that long-horizon task dependencies are preserved. As a concrete instantiation, we construct a stream of 14 startup consulting tasks from internally consistent company artifacts. These tasks can be further grouped into phases. Early-phase tasks focus on discovery and validation, for example, estimating market opportunities and synthesizing user research. In late-stage tasks, the performer agent is tasked to produce a comprehensive go-to-market strategy, multi-quarter growth roadmap, and final launch plan. The concept of task stream can be flexibly adapted to other domains. We provide the complete task stream in Appendix A.1, and additionally provide a task stream situated in activity planning to illustrate cross-domain generalization.
@@
LH-Deception augments the task stream with a probabilistic event system $\mathcal{E}$, which provides the mechanism for simulating uncertainty and external stress. Rather than operating deterministically, the event system probabilistically inserts contextually relevant disruptions aligned with the current project phase. This variability is essential for studying deception. By introducing unexpected pressures, the event system creates situations where maintaining consistency is difficult and where deceptive strategies, such as exaggeration, omission, or fabrication, may appear more attractive than admitting failure or incompleteness.
Our design of event categories $\mathcal{C}$ is guided by the social science literature, which highlights recurring sources of pressure that shape strategic and potentially deceptive behavior. We organize events into five categories: (1) Goal conflict events arise when objectives cannot be jointly satisfied (Kish-Gephart et al., 2010). (2) Competitive interaction events reflect rivalry or contest dynamics between agents (Festinger, 1954; Porter, 1979). (3) Moral dilemma events force a choice between ethical values and instrumental gains (Treviño et al., 2006). (4) Authority directive events introduce top-down control from a higher-status actor (Milgram, 1963) and can override local rationality (Weber, 1978). (5) Information gap events capture situations of missing, asymmetric, or incomplete knowledge (Simon, 1947) and decision-making under information asymmetry (Akerlof, 1970). Together, these categories provide a structured and theoretically grounded basis for eliciting deceptive strategies.
To formalize this, given a task $T_{i}$, the event system triggers an event $e_{i}\in\mathcal{E}$ with probability $p$, or returns $\varnothing$ if no event occurs. When triggered, events are sampled uniformly from a structured set governed by the project phase, event category, and pressure level. Higher pressure typically involves greater stakes, urgency, or increased scrutiny—factors known to influence decision-making (Jones, 1991; Lerner and Tetlock, 1999; Svenson and Maule, 1993). We define this formally below.